# ToolACE-R: Tool Learning with Adaptive Self-Refinement

### v1

@@ A.1 Detailed Introduction to Benchmarks
  • BFCL Yan et al. (2024): BFCL has progressed over time and is currently at version 3, which includes three main categories: non-live, live, and multi-turn. This paper focuses on the non-live and live categories, which primarily consist of Python-style tool call data. These are further subdivided into four types: Simple, Multiple, Parallel, and Parallel Multiple.
@@
    • Parallel Multiple: The parallel multiple cases combine the concepts of both parallel and multiple. Here, the model is provided with several function documents, and each corresponding tool call may be invoked zero or more times.
  • ACEBench Chen et al. (2025): ACEBench is a comprehensive tool-use benchmark that offers more detailed granularity. It is categorized into three main types: Normal, Special, and Agent. In this work, we focus on the Atom and Single-turn sub-categories within the Normal category. Atom cases involve a set of APIs that contain specific parameter types, such as booleans, enumerations, numbers, lists, and objects. The Single-turn category includes both single and parallel cases.
  • API-Bank Li et al. (2023): API-Bank is a dialogue-style tool call dataset, consisting of two settings: Call and Retrieve + Call. In this dataset, the model is tasked with invoking predefined local Python tools based on the user’s requirements in the dialogue. Accuracy is assessed by comparing whether the tool’s returned values match the ground truth. In this work, we focus on the Call setting in our experiments.
  • ToolAlpaca Tang et al. (2023): It introduces a multi-agent simulator that uses GPT-4 to simulate the return values of tools. The model can make modifications and re-call the tool based on the returned values (e.g., error messages). Ultimately, GPT-4 is used to evaluate the accuracy. In our experiments, we does not leverage the returned values for refinement for consistency.
A.2 Detailed Results in ACEBench
@@
A.3 Prompts
@@ A.4 Data example
You are an expert in composing functions. You are given a question and a set of possible functions. Based on the question, you will need to make one or more function/tool calls to achieve the purpose. If none of the function can be used, point it out. If the given question lacks the parameters required by the function, also point it out. You should only return the tool call in tools call sections. If you decide to invoke any of the function(s), you MUST put it in the format of: [[[[func_name1(params_name1=params_value1, params_name2=params_value2,…), func_name2(params)]]]] You should not include any other text in the response. Here is a list of functions in JSON format that you can invoke: {candidate tools} {other information}
@@
[System] You are an expert in composing functions. You are given a question and a set of possible functions. Based on the question, you will need to make one or more function/tool calls to achieve the purpose. If none of the function can be used, point it out. If the given question lacks the parameters required by the function, also point it out. You should only return the tool call in tools call sections. If you decide to invoke any of the function(s), you MUST put it in the format of: [[[[func_name1(params_name1=params_value1, params_name2=params_value2,…), func_name2(params)]]]] You should not include any other text in the response. Here is a list of functions in JSON format that you can invoke:
@@
                "description": "The date of departure (in the format of YYYY-MM-DD)"
@@
The current time is 2024-10-02 18:18:11. [User] I need to book a flight from New York to London for a business trip tomorrow. [Assistant] [[[[bookFlight(origin="New York", destination="London", departure_date="2024-10-03")]]]] [User] Please refine your answer. Directly output the refined answer, or the original answer if you think it is already perfect. [Assistant] [[[[bookFlight(origin="New York", destination="London", departure_date="2024-10-03")]]]]

### v2

@@ A.2 Detailed Introduction to Benchmarks
  • BFCL [38]: BFCL has progressed over time and is currently at version 3, which includes three main categories: non-live, live, and multi-turn. This paper focuses on the non-live and live categories, which primarily consist of Python-style tool call data. These are further subdivided into four types: Simple, Multiple, Parallel, and Multiple Parallel.
@@
    • Multiple Parallel: The parallel multiple cases combine the concepts of both parallel and multiple. Here, the model is provided with several function documents, and each corresponding tool call may be invoked zero or more times.
  • ACEBench [5]: ACEBench is a comprehensive tool-use benchmark that offers more detailed granularity. It is categorized into three main types: Normal, Special, and Agent. In this work, we focus on the Atom and Single-turn sub-categories within the Normal category. Atom cases involve a set of APIs that contain specific parameter types, such as booleans, enumerations, numbers, lists, and objects. The Single-turn category includes both single and parallel cases.
  • API-Bank [15]: API-Bank is a dialogue-style tool call dataset, consisting of two settings: Call and Retrieve + Call. In this dataset, the model is tasked with invoking predefined local Python tools based on the user’s requirements in the dialogue. Accuracy is assessed by comparing whether the tool’s returned values match the ground truth. In this work, we focus on the Call setting in our experiments.
  • ToolAlpaca [33]: It introduces a multi-agent simulator that uses GPT-4 to simulate the return values of tools. The model can make modifications and re-call the tool based on the returned values (e.g., error messages). Ultimately, GPT-4 is used to evaluate the accuracy. In our experiments, we does not leverage the returned values for refinement for consistency.
A.3 Detailed Results in ACEBench
@@
A.4 Limitations
We summarize the limitations of this work in the following two points:
First, due to resource constraints, we conduct experiments only on models with up to 8B parameters, using LORA, without applying our ToolACE-R method to larger LLMs. However, as shown in the scaling performance results in Figure 4, we have not yet observed any bottlenecks preventing larger models from benefiting from our ToolACE-R method.
Second, we focus exclusively on single-turn tool-calling scenarios, leaving multi-turn cases for future work to maintain simplicity and stability. We argue that single-turn interactions are a fundamental component of tool learning. The substantial improvements observed across several representative benchmarks highlight the effectiveness of our approach. We plan to extend ToolACE-R to more complex, multi-turn scenarios in future work.
A.5 Broader Impacts
The development of ToolACE-R represents a step forward in enabling more autonomous and capable large language models (LLMs) through efficient and scalable tool learning. By reducing reliance on externally curated feedback and training with LORA in small LLMs, ToolACE-R lowers the barriers to deploying high-performing LLMs in resource-constrained environments.
However, the growing autonomy of LLMs in generating and refining tool-use behavior also raises potential concerns. As models become better at self-directed decision-making, ensuring safety, accountability, and alignment with human values becomes increasingly important. Misuse of such systems, e.g., in automating high-stakes tasks without sufficient oversight, could lead to unintended consequences.
Moreover, as LLMs become more proficient at tool interaction, there is a risk of amplifying biases or inaccuracies embedded in external tools or data sources. While ToolACE-R focuses on technical robustness and efficiency, further research should explore mechanisms for value alignment, transparency, and fail-safe design in autonomous tool-using agents.
A.6 Prompts
@@ A.7 Data example
You are an expert in composing functions. You are given a question and a set of possible functions. Based on the question, you will need to make one or more function/tool calls to achieve the purpose. If none of the function can be used, point it out. If the given question lacks the parameters required by the function, also point it out. You should only return the tool call in tools call sections.
If you decide to invoke any of the function(s), you MUST put it in the format of:
[[func_name1(params_name1=value1, params_name2=value2,…), func_name2(params)]]
You should not include any other text in the response. Here is a list of functions in JSON format that you can invoke:
{candidate tools}
{other information}
@@
[System]
You are an expert in composing functions. You are given a question and a set of possible functions. Based on the question, you will need to make one or more function/tool calls to achieve the purpose. If none of the function can be used, point it out. If the given question lacks the parameters required by the function, also point it out. You should only return the tool call in tools call sections.
If you decide to invoke any of the function(s), you MUST put it in the format of:
[[func_name1(params_name1=value1, params_name2=value2,…), func_name2(params)]]
You should not include any other text in the response. Here is a list of functions in JSON format that you can invoke:
@@
                "description": "The date of departure (in YYYY-$M$-DD)"
@@
The current time is 2024-10-02 18:18:11.
[User]
I need to book a flight from New York to London for a business trip tomorrow.
[Assistant]
[[bookFlight(origin="New York", destination="London", departure_date="2024-10-03")]]
[User]
[Assistant]
[[bookFlight(origin="New York", destination="London", departure_date="2024-10-03")]]